The sum of jack and his father's age is 52. Jack's father age is 2 less than 5 times jack's age. Find the ages of jack and his father.
Answer by ewatrrr(24785)


Hi,
Let x and (52-x)represent jack's and the father's age respectively
Write as we Read***
(52-x) = 5x - 2
Solving for x
52 - x = 5x - 2
54 = 6x
x = 9, jack's age. The father's age is 43. (52-9)
CHECKING our Answer***
43 = 5*9 - 2